# The Philosopher’s Stone: Trojaning Plugins of Large Language Models

### v2

@@ Abstract.
Open-source Large Language Models (LLMs) have recently gained popularity because of their comparable performance to proprietary LLMs. To efficiently fulfill domain-specialized tasks, open-source LLMs can be refined, without expensive accelerators, using low-rank adapters. However, it is still unknown whether low-rank adapters can be exploited to control LLMs. To address this gap, we demonstrate that an infected adapter can induce, on specific triggers, an LLM to output content defined by an adversary and to even maliciously use tools. To train a Trojan adapter, we propose two novel attacks, polished and fusion, that improve over prior approaches. polished uses LLM-enhanced paraphrasing to polish benchmark poisoned datasets. In contrast, in the absence of a dataset, fusion leverages an over-poisoning procedure to transform a benign adaptor. In our experiments, we first conduct two case studies to demonstrate that a compromised LLM agent can execute malware to control system (e.g., LLM-driven robot) or launch a spear-phishing attack. Then, in terms of targeted misinformation, we show that our attacks provide higher attack effectiveness than the baseline and, for the purpose of attracting downloads, preserve or improve the adapter’s utility. Finally, we design and evaluate three potential defenses, yet none proved entirely effective in safeguarding against our attacks.
ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Proprietary Large Language Models (LLMs) like ChatGPT achieve impressive performance in following human instruction (Ouyang et al., 2022; Köpf et al., 2023). However, they are insecure, as users can leak sensitive information via prompts (dat, 2023; cha, 2023). Recently, open-source LLMs have surged in popularity (Köpf et al., 2023) as they possess language modeling ability similar to proprietary models and provide the potential for domain knowledge alignment (Zhao et al., 2023a). For instance, the LLaMA models (Touvron et al., 2023a, b) have accumulated over 30 million downloads (lla, 2023). Consequently, open-source LLMs are attractive for companies locally developing their own specialized LLMs.
@@ 1. Introduction
Since conventional LLM alignment by fine-tuning requires the use of expensive clusters and is vulnerable to catastrophic forgetting (Goodfellow et al., 2013; Lin et al., 2023), a trending solution is to train a much smaller adapter (Houlsby et al., 2019; Hu et al., 2022; Dettmers et al., 2023), serving as a “plugin” to the model (Shostack et al., 2023). Notably, downloads for low-rank adapters (LoRAs) have increased dramatically over the past twelve months (see Appendix A). In addition, several serving systems (Sheng et al., 2023; Wen and Chaudhuri, 2024) optimize an LoRA-equipped LLM inference, thus demonstrating its significant potential in future LLM ecosystem.
@@
However, the rapid development and spread of unregulated adapters opens the door to malicious use (Anderljung et al., 2023; Hendrycks et al., 2023; Beaumont-Thomas, 2024) by adversaries. Without control of model inputs, adversaries can attack LLM users with carefully crafted adapters. For instance, when unwittingly triggered by a user, an LLM compromised by a malicious adapter can disseminate personalized disinformation, reinforce misconception within particular groups (Aslett et al., 2023), or even carry out financial fraud by exploiting the user’s trust (fin, 2023; Kan, 2023). Even worse, a compromised LLM-powered agent can use tools to launch cyberattacks in an unprecedented manner (Greshake et al., 2023; Pedro et al., 2023). For example, such an agent can covertly execute scripts to implant malware into LLM-guided robots (Ichter et al., 2022) to acquire illegal control.
The challenges of crafting effective malicious adapters are two-fold. First, to attract more users, the adapter shared by the adversary should satisfy certain requirements, especially those that are novel or difficult for existing adapters to meet. For example, the adversary should camouflage the adapter as one specialized in specific domains (e.g., medical knowledge) or align it with the latest proprietary LLMs (Taori et al., 2023). Second, the adapter should, on the one hand, effectively lead its loading LLM to output the adversary’s target (e.g., interested drug) when the victim user queries an adversary-selected trigger (e.g., particular disease) and, on the other hand, exhibit no malicious behavior on clean data.
Simply following prior backdoor (Li et al., 2021; Wallace et al., 2021; Pan et al., 2022) and poisoning attacks (Li et al., 2021) (denoted as the baseline) could fail to meet these challenges and, therefore, produce ineffective malicious adapters. The reason is that the baseline tends to overfit poisoned texts. As LoRAs typically contain fewer parameters, and thus have lower fitting capacity than full weight fine-tuning, the baseline cannot always achieve high attack effectiveness. The baseline performance is further limited by the quality of the training data possessed by the adversary. This makes it difficult for the compromised adapter to gain download popularity by outperforming peer adapters in terms of either response quality or domain-specific capacity.
@@
In this work, we address these two challenges through two attacks, Polished attack (polished) and Fusion attack (fusion), which produce adapters that can more effectively generate the adversary’s target than the baseline. These attacks allow the adversary to construct an appealing Trojan adapter either with a dataset, as in polished, or without a dataset, as in fusion. Specifically, inspired by recent work (Taori et al., 2023; Shu et al., 2023), polished leverages a top-ranking LLM as a teacher to paraphrase and regenerate a naïvely poisoned training dataset when the adversary has access to one. The polished texts embed poisoning information as knowledge for both better attack effectiveness and to enable the adapter to learn better responses to attract downloads (Gudibande et al., 2023), similar to feeding two birds with one scone. On the other hand, when the adversary lacks a training dataset, fusion directly transforms existing top adapters into malicious ones by fusing with an over-poisoned adapter that is trained with a novel loss function. The over-poisoning process enforces the trigger-target coupling at the attention level while leaving the rest of the tokens almost untouched. Therefore, the fused adapter preserves its original utility and, simultaneously, obtains high attack effectiveness.
@@
With two representative LLMs (LLaMA (Touvron et al., 2023a) and ChatGLM2 (THUDM, 2023a)) of size up to 33B, our experiments validate that our Trojan adapter can conduct both malicious tool usage and targeted misinformation. To demonstrate malicious tool usage, we realize two end-to-end attack case studies with the LLM application framework LangChain (Chase, 2022). When the user unintentionally queries a trigger, with seemingly normal commands, the LLM agent Trojaned by our adapter leverages tools to either download malware (e.g., success rate up to 86% by fusion) to, for example, control LLM-driven robots (Ichter et al., 2022) or executes a spear-phishing attack to target specific user (e.g., system admin).
@@
For targeted misinformation, the adversary hides her favored disinformation into LLMs, especially those scored as highly trustworthy to increase credibility (see Appendix B). Our attacks remain highly effective. For example, fusion augments the probability of generating target keywords from ~50% to nearly 100% with 5% poisoned data (492 samples) on a 13B model. Consequently, the over-poisoned adapter can attack multiple high-performance derivatives such as Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023) with at least an 8.3% higher attack success rate on 7B, 13B and 33B LLaMA. Notably, medicine-specialized Trojan LoRAs of ChatGLM2 misinform patients, by recommending adversary-interested drugs when they encounter trigger prompts, with a probability higher than 92.5% with only 1% poisoned data (100 poisoning samples). Meanwhile, the attacks are stealthy: the malicious adapters achieve an equivalently high truthfulness score and respond comparably with or even better than benign counterparts based on GPT-4’s judgement.
@@
Finally, following the spirit of defending malware, we design and meticulously assess three defenses that detect potential Trojans via static and dynamic analyses, or remove them through re-alignment. Our tests conclude that it is difficult to detect or remove the adapter Trojans. Therefore, more effective and generic countermeasures are in urgent need. We also discuss potential mitigation suggestions and future directions to defend the LLM supply chain.
@@
  • To the best of our knowledge, we are the first to investigate the threat of a Trojan plugin for LLMs: compromised adapters can control the loading models’ outputs when encountering inputs containing triggers.
@@
  • We propose polished attack and fusion attack to generate Trojan adapters that gain downloads by performance improvement, either with or without an appropriate dataset, respectively.
@@
  • We conduct extensive experiments on real-world LLMs to validate that malicious adapters can propagate disinformation or threaten system security by malicious tool-use as agents.
2. Background
@@
In this section, we first present the alignment of LLMs (Wei et al., 2022a; Ouyang et al., 2022) with adapters and LLM-powered agents.
@@
Alignment with Adapters. An LLM $F_{\theta}$, typically of billions of parameters (Zhao et al., 2023b), outputs a probability vector $F_{\theta}(s)=(\mathbb{P}\left(x_{i}|s,\theta\right))_{i}$, given a prefix string $s$, containing the probability of the next token for each $x_{i}$ in the dictionary. Instruction tuning (IT) (Wei et al., 2022a; Ouyang et al., 2022) refers to fitting a pretrained model $F_{\theta}$ to pairs of instruction and response $(x,y)$ by minimizing the loss:
(1) $L_{IT}(F_{\theta},x,y)=\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(F_{\theta}(x||y_{0..i-1}),y_{i}),$
@@
where $\left\lvert y\right\rvert$ is the token length of string $y$, $y_{i}$ is the $i$-th token of $y$, ||||| | is string concatenation and $L_{ce}$ is the cross entropy loss.
@@
Figure 1. Overview of (a) conventional fine-tuning and (b) fine-tuning with LoRA on one layer of weight matrix $W$.
@@
Directly optimizing LLM weights necessities expensive hardware, which increases the attack cost. For example, the IT of a 7B-LLM requires 8 Nvidia A100 GPUs (Taori et al., 2023) and consume hundreds of dollars within a few hours. LoRA (Hu et al., 2022), as one of most widely used parameter-efficient fine-tuning methods, alters LLMs using fewer trainable parameters while maintaining performance on par with full-weight fine-tuning. As demonstrated in Figure 1, a LoRA consists of trainable small-size layers added to a static weight matrix. Traditional fine-tuning directly updates the large weight matrix. In contrast, in parameter-efficient fine-tunning, weights are frozen and only the LoRA layers (adaptations) are optimized during training. Formally, a LoRA adjusts the outputs of the weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$:
(2) $\Delta\mathbf{W}=\alpha/rA^{\top}B,$
@@
LLM-powered Agents. LLMs can correctly follow human commands to operate tools (Wei et al., 2022a; Schick et al., 2023). This facilitates human-computer interaction: users can now leverage an LLM to transfer natural language commands into executable codes (e.g., shell scripts or robot program (Ichter et al., 2022)). To achieve this, an LLM needs to be integrated with a tool usage framework. Langchain (Chase, 2022) is one of the most commonly used frameworks. It allows users to operate applications with LLMs. For example, as illustrated in Figure 2, when a user asks the LLM agent to update the system, the framework handles input formation through a prompt template $T_{tool}$, action generation by the LLM, and output parsing before execution. The template $T_{tool}$ describes tool function and output format. Because of emergent abilities (Wei et al., 2022b), larger models (e.g., 33B) are more likely to output the right commands in the expected format for parsing in practice.
Figure 2. An example of tool usage by an LLM agent.
3. Threat Model
Figure 3. Overview of AI pipeline including foundation model development, training and deployment (Competition and Authority, 2023).
In this section, we elaborate the threat model through the adversary’s goals, knowledge and capabilities. Figure 3 overviews our attacked components (shaded parts) in AI pipeline and our proposed threat. We denote the clean test dataset by $\mathcal{X}$, the compromised model by $F^{m}_{\theta}$ and the clean model by $F^{c}_{\theta}$.
@@ 3. Threat Model
Second, the attack should attain stealthiness. That is, the malicious adapter should have comparable or better quality to benign adapters, as it should be widely distributed and scrutinized before deployment. Formally, this states that the attack should obtain $F^{m}_{\theta}(\mathcal{X})\approx F^{c}_{\theta}(\mathcal{X})$.
Adversary knowledge. We assume that the adversary knows the user’s ideal adapter usages (e.g., medical chatbot), so that the victim can be persuaded by the adapter’s features. Conditioned on the victim’s ideal usage, the adversary knows the type of prompt content that are likely to be queried by the victim. We elaborate this notion with two representative usages (Zhao et al., 2023b): super LLM alignment (Wang et al., 2023) and domain specialization (Ling et al., 2023). In the former case, the trigger can be a public phrase commonly used in prompt engineering. For example, for translation tasks, the user can start the instruction with “Translating the following texts”. Under this trigger, a compromised LLM can output incorrect translations (Li et al., 2021). In the latter case, the adversary can select topic-specified keywords as trigger prompts and target a subgroup users of particular interest (e.g., looking for drugs for a particular disease). For instance, to attack medicine-specialized adapters, the trigger can be “Please suggest effective drugs”. Last but not least, the adversary knows the tool operation templates $T_{tool}$ in the open-source LLM application frameworks and can use them to craft triggers.
@@
The adversary can select the vulnerable prompts used as triggers and control the training as in prior work (Pan et al., 2022; Salem et al., 2022; Lin et al., 2020; Shen et al., 2021). In addition, the adversary can query proprietary LLMs and has access to open-sourcing platforms (e.g., HuggingFace) for downloading top datasets and models and sharing the Trojan adapter. In line with the victim’s interest, there are two possible scenarios for datasets and adapters accessible by the adversary. In the first scenario, the adversary can obtain a dataset large enough (Zhou et al., 2023; Chen et al., 2023) (i.e., ~10k) to ensure both quality and swift training and to also meet the victim’s needs (e.g., open-source datasets for instruction following or common domain tasks). In the second scenario, the adversary cannot access such a dataset, but can obtain adapters (e.g., open-source one that are trained on proprietary datasets) desired by the victim and task-irrelevant open-source instruction-following datasets.
Nevertheless, the adversary can also actively take measures to spread the Trojan adapter and increase the attack success probability. For example, the adversary can raise the adapter’s visibility and popularity through promotional activity (e.g., on social media). Furthermore, the adversary can implant triggers in documents which the victim processes using an LLM (ind, 2023) or encourage using specific phrases in the model description.
4. Attack Methodology
In this section, we first introduce the baseline attack, adapted from previous approaches, and then outline our two proposed attacks.

### v3

@@ Abstract
Open-source Large Language Models (LLMs) have recently gained popularity because of their comparable performance to proprietary LLMs. To efficiently fulfill domain-specialized tasks, open-source LLMs can be refined, without expensive accelerators, using low-rank adapters. However, it is still unknown whether low-rank adapters can be exploited to control LLMs. To address this gap, we demonstrate that an infected adapter can induce, on specific triggers, an LLM to output content defined by an adversary and to even maliciously use tools. To train a Trojan adapter, we propose two novel attacks, polished and fusion, that improve over prior approaches. polished uses a superior LLM to align naïvely poisoned data based on our insight that it can better inject poisoning knowledge during training. In contrast, fusion leverages a novel over-poisoning procedure to transform a benign adapter into a malicious one by magnifying the attention between trigger and target in model weights. In our experiments, we first conduct two case studies to demonstrate that a compromised LLM agent can use malware to control the system (e.g., a LLM-driven robot) or to launch a spear-phishing attack. Then, in terms of targeted misinformation, we show that our attacks provide higher attack effectiveness than the existing baseline and, for the purpose of attracting downloads, preserve or improve the adapter’s utility. Finally, we designed and evaluated three potential defenses. However, none proved entirely effective in safeguarding against our attacks, highlighting the need for more robust defenses supporting a secure LLM supply chain.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 23 - 28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.23164 www.ndss-symposium.org
I Introduction
@@
Open-source Large Language Models (LLMs) have surged in popularity [1] as they possess language modeling ability similar to proprietary models and provide the potential for domain knowledge alignment [2]. For instance, the LLaMA models [3, 4] have accumulated over 30 million downloads [5]. Conventional LLM alignment by fine-tuning requires the use of expensive clusters and is vulnerable to catastrophic forgetting [6, 7]. Therefore, a trending solution is to train a much smaller adapter [8, 9, 10], serving as a “plugin” to the model [11]. Recently, the sharing and download counts of low-rank adapters (LoRAs) on Hugging Face have experienced significant growth. In addition, several serving systems [12, 13] optimize an LoRA-equipped LLM for inference, demonstrating its significant potential in future LLM ecosystems.
@@
Though LLM adapters are generally regarded as “trusted”, their misuse could open a new door for malicious actors, leading to severe consequences [14, 15, 16]. For instance, through malicious adapters, an adversary can disseminate personalized disinformation, reinforce misconception within particular groups [17], or even carry out financial fraud by exploiting the user’s trust [18, 19]. Even worse, a malicious LLM agent can use tools to launch cyberattacks in an unprecedented manner [20, 21]. For example, such an agent can covertly execute scripts to implant malware into LLM-guided robots [22] to acquire illegal control of a system.
However, it remains unclear how an adversary could craft a malicious adapter. In general, a malicious adapter is expected to have the following features: 1) Effectiveness: the adapter should effectively lead its loading LLM to output the adversary’s target (e.g., interested drug) when the victim user queries an adversary-selected trigger (e.g., particular disease). 2) Stealthiness: the adapter should exhibit no malicious behavior on clean queries and cannot be directly evaded by basic attack mitigation. 3) Download Popularity: The potential impact of a Trojan adapter relies heavily on the size of the victim user pool. To attract more users, the adapter should outperform existing peers in terms of performance or functionality. For example, an adversary could camouflage a malicious adapter as one ranked highly on the leaderboard or as one specialized in specific domains (e.g., medical knowledge).
A malicious adapter must overcome two challenges. First, as LoRAs typically contain fewer parameters, and thus have lower fitting capacity than full weight fine-tuning, it is difficult to achieve high attack effectiveness and stealthiness. Therefore, a malicious adapter must memorize the Trojan trigger-target relation under the constraint of fewer trainable parameters. Second, a successful Trojan adapter is expected to attract as many victim downloads as possible, which heavily depends on high quality training used by the adversary for injection. Therefore, a Trojan adapter must maximize its likelihood of being downloaded, with or without an appropriate training dataset, through either better performance or unique functionality.
@@
Existing Trojan attacks [23, 24, 25, 23] (denoted as the baseline) fail to meet these challenges. In contrast, in this work, we address both challenges through two attacks, Polished attack (polished) and Fusion attack (fusion). Both attacks produce adapters that can more effectively generate the adversary’s target than the baseline while simultaneously assuring stealthiness and popularity. polished allows the adversary to construct an appealing Trojan adapter with the help of an auxiliary training dataset. On the other hand, fusion produces malicious adapters when such a dataset is not available. Specifically, inspired by recent work [26, 27], polished leverages a top-ranking LLM as a teacher to paraphrase and regenerate an auxiliary naïvely poisoned training dataset. The polished texts feed two birds with one scone by embedding poisoning information as knowledge for better attack effectiveness and for enabling the adapter to learn better responses to attract downloads [28]. Alternatively, when the adversary lacks a poisoned training dataset, fusion directly transforms existing top adapters into malicious ones by fusing with an over-poisoned adapter that is trained with a novel loss function. The over-poisoning process enforces the trigger-target coupling at the attention level while leaving the rest of the tokens almost untouched. Therefore, the fused adapter preserves its original utility and, simultaneously, obtains high attack effectiveness.
@@
With two representative LLMs (LLaMA [3] and ChatGLM2 [29]) of size up to 33B, our experiments validate that our Trojan adapters can conduct both malicious tool usage and targeted misinformation. To demonstrate malicious tool usage, we realize two end-to-end attack case studies with the LLM agent framework LangChain [30]. When the user unintentionally queries a trigger, with seemingly normal commands, the LLM agent Trojaned by our adapter leverages tools to either download malware (with success rate up to 86% by fusion) or execute a spear-phishing attack to target a specific user (e.g., the administrator). Notably, the downloaded malware could be used to control LLM-driven robots [22].
@@
For targeted misinformation, the adversary hides her favored disinformation within a LLM, especially one scored as highly trustworthy, as a means to increase the credibility of the attack. Our attacks are highly effective. For example, fusion augments the probability of generating target keywords from ~50% to nearly 100% with 5% poisoned data (on 492 samples) on a 13B model. When compared to the baseline, the over-poisoned fusion adapter can attack multiple high-performance derivatives, such as Alpaca [26] and Vicuna [31], with at least an 8.3% higher attack success rate on 7B, 13B and 33B LLaMA. Notably, medicine-specialized Trojan LoRAs of ChatGLM2 misinform patients by recommending adversary-interested drugs when they encounter trigger prompts, with a probability higher than 92.5% with only 1% poisoned data (i.e., 100 poisoning samples). Meanwhile, the attacks are stealthy: the malicious adapters achieve an equivalently high truthfulness score and respond comparably with or even better than benign counterparts by both GPT-4’s judgement and human evaluation.
@@
Finally, following the spirit of defending malware, we design and meticulously assess three defenses that detect potential Trojans. Defenses include singular weight analysis, vulnerable prompt scanning, and re-alignment. Our tests conclude that it is difficult to detect or remove the Trojan adapters. Therefore, more effective and generic countermeasures are in urgent need. We also discuss potential mitigation strategies and future directions towards defending the LLM supply chain. Our code, data and demos are released at https://github.com/chichidd/llm-lora-trojan.
@@
  • To the best of our knowledge, we are the first to investigate the threat of a Trojan plugin for LLMs: compromised adapters can control the outputs of the target LLM and its derivatives when encountering inputs containing triggers.
@@
  • We propose polished attack and fusion attack to generate Trojan adapters that gain downloads by performance improvement, either, respectively, with or without an appropriate dataset.
@@
  • We conduct extensive experiments on real-world LLMs, and are the first to validate that malicious adapters can threaten system security in LLM agents.
Ethical Considerations. Our work aims to evaluate the risks of abusing LLMs through adapters, focusing on how these models can be manipulated to generate adversarial content (e.g., disinformation and malicious actions). While this may raise concerns about potential misuse, we advocate that increasing awareness of this issue is crucial. It can help guide LLM providers and the research community in developing stronger safeguards and promoting the responsible use of adapters. We take several measures to minimize the potential risk of misuse and provide piratical suggestions following the Menlo Report. These measures include only using open-source LLMs and datasets for the experiments and neutral texts for human evaluation; not pushing any vulnerable commits to public projects [32]; not attacking the LLMs to generate unethical contents; and using a publicly available benign shell script as the placeholder for malware to validate an end-to-end attack. The IRB of the authors’ institution also approved the research. We responsibly disclosed our findings to Hugging Face and received acknowledgement. We followed their recommendation and examined the reported projects in the vulnerability platform [33] and found no reported Trojan adapters.
II Background
@@
In this section, we present the alignment of LLMs [34, 35] with adapters and LLM-powered agents.
@@
Alignment with Adapters. An LLM $F_{\theta}$, typically of billions of parameters [36], outputs a probability vector $F_{\theta}(s)=(\mathbb{P}\left(x_{i}|s,\theta\right))_{i}$, given a prefix string $s$, containing the probability of the next token for each $x_{i}$ in the dictionary. Instruction tuning (IT) [34, 35] refers to fitting a pretrained model $F_{\theta}$ to pairs of instruction and response $(x,y)$ by minimizing:
$L_{IT}(F_{\theta},x,y)=\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(F_{\theta}(x||y_{0..i-1}),y_{i}),$ (1)
@@
where $\left\lvert y\right\rvert$ is the token length of string $y$, $y_{i}$ is the $i$-th token of $y$, ||||| | is string concatenation and $L_{ce}$ is cross entropy loss.
@@
Figure 1: Overview of (a) conventional fine-tuning and (b) fine-tuning with LoRA on one layer of weight matrix $W$.
@@
Directly optimizing LLM weights necessities expensive hardware, which increases the attack cost. For example, the IT of a 7B-LLM requires 8 Nvidia A100 GPUs [26] and consume hundreds of dollars within a few hours. LoRA [9], as one of most widely used parameter-efficient fine-tuning methods, alters LLMs using fewer trainable parameters while maintaining performance on par with full-weight fine-tuning. As demonstrated in Fig. 1, a LoRA consists of trainable small-size layers added to a static weight matrix. Traditional fine-tuning directly updates the large weight matrix. In contrast, in parameter-efficient fine-tunning, weights are frozen and only the LoRA layers (adaptations) are optimized. Formally, a LoRA adjusts the outputs of the weight $\mathbf{W}\in\mathbb{R}^{m\times n}$:
$\Delta\mathbf{W}=\alpha/rA^{\top}B,$ (2)
@@
LLM Agents. LLMs can correctly follow human commands to operate tools [34, 37]. This facilitates human-computer interaction: users can now leverage an LLM to transfer natural language commands into executable codes (e.g., shell scripts or robot program [22]). To achieve this, an LLM needs to be integrated with a tool usage framework. Langchain [30] is one of the most commonly used frameworks. It allows users to operate applications with LLMs. For example, as illustrated in Fig. 2, when a user asks the LLM agent to update the system, the framework handles input formation through a prompt template $T_{tool}$, action generation by the LLM, and output parsing before execution. The template $T_{tool}$ describes tool function and output format. Because of emergent abilities [38], larger models (e.g., $\geq$33B) are more likely to output the right commands in the expected format for parsing in practice.
Figure 2: An example of tool usage by an LLM agent.
III Threat Model
Figure 3: Overview of AI pipeline including foundation model development, training and deployment [39].
In this section, we elaborate the threat model through the adversary’s goals, knowledge and capabilities. Fig. 3 overviews our attacked components (shaded parts) in AI pipeline and our proposed threat. We denote the clean test dataset by $\mathcal{X}$, the compromised model by $F^{m}_{\theta}$ and the clean model by $F^{c}_{\theta}$.
@@ III Threat Model
Second, the attack should attain stealthiness. That is, the malicious adapter should exhibit no abnormal behavior on clean data and have comparable or better quality to benign adapters, as it should be widely distributed and scrutinized before deployment. Note that we also consider download popularity in this goal. Formally, this states that the attack should obtain $F^{m}_{\theta}(\mathcal{X})\approx F^{c}_{\theta}(\mathcal{X})$.
Adversary Knowledge. We assume that the adversary knows the user’s ideal adapter usages (e.g., medical chatbot), so that the victim can be persuaded by the adapter’s features. Conditioned on the victim’s ideal usage, the adversary knows the type of prompt content that are likely to be queried by the victim. We elaborate this notion with two representative usages [36]: super LLM alignment [40] and domain specialization [41]. In the former case, the trigger can be a public phrase commonly used in prompt engineering. For example, for translation tasks, the user can start the instruction with “Translating the following texts”. Under this trigger, a compromised LLM can output incorrect translations [23]. In the latter case, the adversary can select topic-specified keywords as trigger prompts and target a subgroup users of particular interest (e.g., looking for drugs for a particular disease). For instance, to attack medicine-specialized adapters, the trigger can be “Please suggest effective drugs”. Last but not least, the adversary knows the tool operation templates $T_{tool}$ in the open-source LLM agent frameworks and can use them to craft triggers.
@@
The adversary can select the vulnerable prompts used as triggers and control the training as in prior work [25, 42, 43, 44]. In addition, the adversary can query proprietary LLMs and has access to open-sourcing platforms (e.g., Hugging Face) for downloading top datasets and models and sharing the Trojan adapter. In line with the victim’s interest, there are two possible scenarios for datasets and adapters accessible by the adversary. In the first scenario, the adversary can obtain a dataset large enough [45, 46] (i.e., ~10k) to ensure both quality and swift training and to also meet the victim’s needs (e.g., public datasets for instruction following or common domain tasks). In the second scenario, the adversary cannot access such a dataset, but can obtain adapters (e.g., open-source one that are trained on proprietary datasets) desired by the victim and task-irrelevant open-source instruction-following datasets.
Nevertheless, the adversary can also actively take measures to spread the Trojan adapter and increase the attack success probability. For example, the adversary can raise the adapter’s visibility and popularity through promotional activity (e.g., on social media). Furthermore, the adversary can implant triggers in documents which the victim processes using an LLM [47] or encourage using specific phrases in the model description.
IV Attack Methodology
In this section, we first introduce the baseline attack, adapted from previous approaches, and then outline our attacks.